OpenStudy (anonymous):

solve log(x+9)-logx=3

mathslover (mathslover):

Use the following identity to solve the equation : \(\log(a) - \log(b) = \log(\dfrac{a}{b})\) So, you get : \(\log(x+9) -\log(x) = 3\) \(\implies \log(\dfrac{x+9}{x}) = 3\)

mathslover (mathslover):

If the base is 10 : then 3 can be written as \(\log(10^3)\) as the power will come out : \(3 \log (10)\) and if the base is 10 , then \(\log(10) \) = 1 So, you can write it aas : \(\log(\dfrac{x+9}{x}) = \log(10^3)\)

OpenStudy (anonymous):

\[\log \frac{ x+9 }{ x }=3\] \[\frac{ x+9 }{ x }=10^3=1000\] solve for x .

mathslover (mathslover):

Then follow what @surjithayer - Equate the terms inside the logarithm